OpenStudy (anonymous):

help me plzzzzzz 1. Solve 5x = 15 for x. 1 2 3 4 2. Solve -11x = 22 for x. -1 -2 1 2 3. Solve 3x - 4 = 11 for x. 2 3 4 5 4. Solve 3x + 7 = -5 for x. -4 -2 2 4 5. Solve -3x + 2 = 23 for x. 3 7 -3 -7

jimthompson5910 (jim_thompson5910):

please only post one question at a time

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started 5x = 15 5x/5 = 15/5 ... Divide both sides by 5. 1x/1 = 3/1 x = 3

OpenStudy (ja1):

my my way to many questions for one post, don't piggyback them do them one by one we will be glad to help you :)